OpenStudy (anonymous):

What is the area of a triangle whose base is 12 m and whose height is 18 m? A.60 B. 108 C. 216 D. 308

OpenStudy (anonymous):

12*18=216/2=108 your answer is 108 so B

OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 11.5 yd and a width of 5 yd? (Points : 1) 66 yd 33 yd 18 yd 16.5 yd

OpenStudy (anonymous):

11.5*2 + 5*2 23+10 33

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

12*18=216 216 divided by 2=108 B.108

OpenStudy (anonymous):

Thank you guys I got a 100%

OpenStudy (anonymous):

S = 12*18/2=216/2=108 ... so B... 108